OpenStudy (anonymous):

What is the equation in standard form of the line which passes through (-2, -3) and has a slope of 1/2 2x + y = 4 2x - y = 4 x + 2y = 4 x - 2y = 4

OpenStudy (cwrw238):

use the standard form of a straight line y - y1 = m(x - x1) where m = slope and (x1,y1) is a point on the line and rearrange to get one of the above

OpenStudy (anonymous):

You can take the points and put them in to the equation to see which one would be right. hint include the negtive sign too

OpenStudy (anonymous):

I will give you an example and work one out. 2x + y = 4 2(-2)+(-3)=4 -4+(-3)=4 -7=4 so this one is wrong do this for all of them.